OpenStudy (anonymous):

Evaluate 2x − y − z when x = −2, y = 3, and z = −2

OpenStudy (anonymous):

Substitute.\[2(-2)-(3)-(-2)=-4-3+2=\]

OpenStudy (anonymous):

with this one you want to put the variables in place for example 2 ( -2) - 3 do you see what I started

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can you show me the finished problem by placing them in

OpenStudy (anonymous):

ireally don't know how to do that

OpenStudy (anonymous):

okay \[2(-2) - 3 - (-2)\]

OpenStudy (anonymous):

that is what it will look like now where do you think we should start

OpenStudy (anonymous):

2-2

OpenStudy (anonymous):

that is 2 * -2

OpenStudy (anonymous):

so 2 * -2 = -4

OpenStudy (anonymous):

so then we have -4 - 3 - (-2)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

but you can change the end into and addition problem because of the two negative signs do that would make it -4 - 3 + 2

OpenStudy (anonymous):

so what is -4 -3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah what?

OpenStudy (anonymous):

what is -4 - 3 =

OpenStudy (anonymous):

-7